OpenStudy (anonymous):

The minimum frequency of light needed to eject electrons from a metal is called the threshold frequency, ν0. Find the minimum energy needed to eject electrons from a metal with a threshold frequency of 4.88 × 1014 s–1. With what kinetic energy will electrons be ejected when this metal is exposed to light with a wavelength of λ = 255 nm?

OpenStudy (anonymous):

I need help with how you do this

OpenStudy (aaronq):

Find the "excess" energy the photon provides to the electron: \(KE=\dfrac{hc}{\lambda}-\phi_0\) The energy of the threshold, \(\phi_0\), is found with the frequency, \(\nu\): \(KE=\dfrac{hc}{\lambda}-h\nu=h(\dfrac{c}{\lambda}-\nu)\)